OpenStudy (anonymous):

A number, twice that number, and one-third of that number are added. The result is 20. What is the original number?

OpenStudy (anonymous):

Since you only have enough information to write one equation, you should write everything in terms of one variable n + 2n + 1/3n = 20 3n + 1/3n = 20 10/3 n = 20 n = 6 now lets check our work: 6 + 12 + 2 = 20 so it is true.